### v2

@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
Theorem 5 clarifies the nature of Conformal Prediction: Although often billed as an uncertainty quantification tool, CP furnishes a representation of predictive uncertainty, but not a quantification in the cardinal sense. The distinction is subtle yet essential: The diameter $\text{diam}(\mathscr{R})$ of a Conformal Prediction Region (CPR) $\mathscr{R}\subseteq Y$ is only defined up to strictly order-preserving (monotone) transformations. [^7] Having assumed a metric $d_{Y}$ on $Y$, $\text{diam}(A)=\sup_{y_{1},y_{2}\in A}d_{Y}(y_{1},y_{2})$, $A\subseteq Y$. A similar argument holds if we consider the volume in place of the diameter of the CPR. Therefore,
[…]
Conformal regions support ordinal comparisons (“larger” versus “smaller”) but not the cardinal arithmetic needed for expected-length-type calibration that variance-based approaches provide. Interpreting region width as a direct proxy for probability mass is therefore as misleading as treating a ranking as if it were a probability distribution.
The commuting diagram, however, reveals a hidden layer: Computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ lives in the space of probability measures, existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024); Chau et al., (2025)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
@@
    Dropping (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with its usual, non-compact topology, take $n=1$, and define $\psi(y^{1},y)\equiv-y$ (jointly continuous, so (ii) stays true). Fix the closed set $A\!\coloneqq(-\infty,0]$. For $y^{1}=(x)$ a short calculation gives
    $\overline{\Pi}_{x}(A)=\begin{cases}1,&x\leq 0,\\[2.0pt] \tfrac{1}{2},&x>0.\end{cases}$
    Hence $x\mapsto\overline{\Pi}_{x}(A)$ has a downward jump at $x=0$, so the envelope is not u.s.c. Consequently $\mathsf{CRED}$ (and therefore $\mathsf{IHDR}$) is not u.h.c. Compactness of $Y$ is thus indispensable.
@@
    Dropping (ii) (joint continuity of $\psi$). Let $Y=[0,1]$ (compact Hausdorff) and $n=1$. Define the permutation-invariant but nowhere continuous score
    $\psi(y^{1},y)=\begin{cases}1,&y\in\mathbb{Q},\\ 0,&y\notin\mathbb{Q}.\end{cases}$
    Fix $\alpha=\tfrac{3}{4}$ and $y^{1}=(0)$. Then $\pi(y,y^{1})=1\text{ if }y\notin\mathbb{Q},$ and $\pi(y,y^{1})=\tfrac{1}{2}\text{ if }y\in\mathbb{Q}.$ Hence $\kappa(\alpha,y^{1},\psi)=[0,1]\cap\mathbb{Q}^{c}$, which is dense but not closed. Because nearby samples $y^{1}=(x)$ with $x\downarrow 0$ randomly interchange rationals and irrationals, the graph of $\kappa$ fails to be closed at $(\alpha,y^{1},\psi)$, so $\kappa$ is not u.h.c. Joint continuity of $\psi$ is therefore necessary.
@@
    Dropping (iii) (no-tie rule). Work with any continuous $\psi\in\mathscr{F}$ (e.g. the neural $\psi_{\theta}$ of Example 1) and fix $n\geq 1$. Choose $\alpha=1\in S_{n+1}$ (which violates the no-tie exclusion). If all $n$ observed points coincide, then every candidate $y$ ties with the sample, giving $\pi(y,y^{n})=1$ for all $y$. Because $\kappa$ uses the strict inequality $\pi>\alpha$, we obtain $\kappa(1,y^{n},\psi)=\emptyset$. For the sequence $1-\tfrac{1}{k}\eqqcolon\alpha_{k}\downarrow 1$ we have $\kappa(\alpha_{k},y^{n},\psi)=Y$ for every $k$, so $\kappa$ is not u.h.c. at $\alpha=1$. The no-tie parameter restriction is thus essential.
@@
    Dropping (iv) (topologies on $\Delta_{Y}$ and $\mathscr{C}$).
    1. (a)
      Replace the weak topology on $\Delta_{Y}$ by the total-variation (TV) norm. Although individual evaluations $P\mapsto P(A)$ remain 1-Lipschitz in TV, $\Delta_{Y}$ is no longer compact, so the Closed Graph Theorem (Lemma 3) breaks down. A concrete example: Take $Y=[0,1]$, fix any $(y^{n},\psi)$, and consider the sequence of Dirac measures $P_{k}\coloneqq\delta_{1/k}\in\mathsf{CRED}(y^{n},\psi)$. The graph-points $\bigl{(}(y^{n},\psi),P_{k}\bigr{)}$ have no TV-convergent subsequence in $(Y^{n}\times\mathscr{F})\times\Delta_{Y}$, so $\mathsf{CRED}$ fails to be u.h.c. Compactness of $\Delta_{Y}$ in the weak topology is therefore indispensable.
    2. (b)
      Keep the weak topology on $\Delta_{Y}$ but equip $\mathscr{C}$ with the Hausdorff metric instead of the Vietoris topology. Let $Y=\{0,1\}$ so that $\Delta_{Y}=[0,1]$ via $p\mapsto P(\{1\})$, and write $A\coloneqq\{1\}$. Define $\mathcal{M}_{k}\coloneqq[0,1-\tfrac{1}{k}]$ and $\mathcal{M}_{\infty}\coloneqq\{1\}$. Then $d_{H}(\mathcal{M}_{k},\mathcal{M}_{\infty})=\tfrac{1}{k}\to 0$, yet $\underline{P}_{\!\mathcal{M}_{k}}(A)=0\to 0<1=\underline{P}_{\!\mathcal{M}_{\infty}}(A),$ so $\underline{P}_{\mathcal{M}}(A)$ is not lower semicontinuous, and $\mathsf{IHDR}$ is not u.h.c. The Vietoris topology on $\mathscr{C}$ is therefore indispensable.
@@
An in-depth study of some of the properties of $\mathbf{UHCont}$ is carried out in Section 7. In the next section, instead, we show that a result similar to Theorem 5 can be derived also for a category whose morphisms are weakly measurable correspondences.
4. Conformal Prediction as a Weakly measurable Morphism
This section introduces category $\mathbf{WMeas}_{\text{uc}}$, and shows that, under minimal assumptions, the (Full) Conformal Prediction correspondence $\kappa$ is a morphism of $\mathbf{WMeas}_{\text{uc}}$ as well, embedded in a commuting diagram. The following too is inspired by Perrone, (2024, Example 1.1.14).
@@ 4. Conformal Prediction as a Weakly measurable Morphism
  • Objects: Measurable Polish spaces $(X,\Sigma_{X})$, with $\Sigma_{X}=\mathcal{B}(X)$, the Borel $\sigma$-algebra.
@@
  • Morphisms: Weakly measurable (see Definition 27 in Appendix A), uniformly compact-valued (see Definition 28 in Appendix A) correspondences $\Phi:X\rightrightarrows Y$.
Theorem 7 ($\mathbf{WMeas}_{\text{uc}}$ is a category).
The structure $\mathbf{WMeas}_{\text{uc}}$ defines a category.
Proof.
We verify the four axioms.
(1) and (4) Identity morphisms and Identity laws. For any object $X$, $\operatorname{id}_{X}(x)=\{x\}$ is non-empty compact. Put $K_{\operatorname{id}_{X}}=X$ (compact because $X$ is Polish and we may take the closed unit ball in its metric); The uniformly compact-valued condition holds. For every open $O\subseteq X$ we have $\operatorname{id}_{X}^{-1}(O)=O\in\Sigma_{X}$, so $\operatorname{id}_{X}$ is a morphism. The usual set-theoretic equalities $\Phi\circ\operatorname{id}_{X}=\Phi$ and $\operatorname{id}_{Y}\circ\Phi=\Phi$ prove the identity laws.
Compactness of fibers. For each $x\in X$,
Since $K_{\Psi}$ is compact and the union is taken inside it, $(\Psi\circ\Phi)(x)$ is a closed subset of a compact space, hence compact. Thus compactness holds with the same compact bound $K_{\Psi}$, so the uniformly compact-valued condition also holds.
Weak measurability. Fix an open $O\subseteq Z$ and define $H\coloneqq\Psi\circ\Phi$. Because $\Phi$ has non-empty closed values, the Kuratowski-Ryll-Nardzewski theorem (Aliprantis and Border,, 2006, Theorem 18.13) yields Borel selectors $(f_{n})_{n\in\mathbb{N}}\subseteq\mathcal{B}(X,Y)$ with $f_{n}(x)\in\Phi(x)$ and $\overline{\{f_{n}(x):n\in\mathbb{N}\}}=\Phi(x)$. Then
$H(x)\cap O\neq\emptyset\iff\exists n\in\mathbb{N}:\Psi(f_{n}(x))\cap O\neq\emptyset,$
so
$H^{-1}(O)=\bigcup_{n\in\mathbb{N}}f_{n}^{-1}\!\bigl{(}\Psi^{-1}(O)\bigr{)}.$
Each $f_{n}$ is Borel and $\Psi$ is weakly measurable, whence $\Psi^{-1}(O)\in\Sigma_{Y}$ and $f_{n}^{-1}\!\bigl{(}\Psi^{-1}(O)\bigr{)}\in\Sigma_{X}$. A countable union of Borel sets is Borel, so $H^{-1}(O)\in\Sigma_{X}$. Therefore $H$ is weakly measurable, completing closure.
Therefore, $\mathbf{WMeas}_{\text{uc}}$ satisfies the axioms of a category. ∎
Similarly to what we did in Section 3, we now ask ourselves what are the minimal conditions that make correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ uniformly compact-valued and weakly measurable. As we shall see in Theorem 8, they are the same as in Section 3, with (i) and (ii) slightly strengthened to
  1. (i’)
    $Y$ is compact metric and $\Sigma_{Y}$ is its Borel $\sigma$-field.
  2. (ii’)
    Each $\psi\in\mathscr{F}$ is jointly continuous on $Y^{n}\times Y$; $\mathscr{F}$ is Polish in the uniform ($\sup$-norm) topology.
Assume that (i’), (ii’), (iii), (iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram
commutes in $\mathbf{WMeas}_{\text{uc}}$.
Proof of Theorem 8.
Uniformly compact-valued. All three images lie inside the fixed compact spaces $Y$ (by (i’)) or $\Delta_{Y}$ (by (iv)); We are now left to show that they themselves are compact-valued.
An argument similar to point (i) in the Proof of Theorem 5 shows that, for all $y^{n}\in Y^{n}$, all $\psi\in\mathscr{F}$, and all $\alpha\in[0,1]\setminus S_{n+1}$, the set $\kappa(\alpha,y^{n},\psi)=\{y:\pi(y,y^{n},\psi)>\alpha\}=\{y:\pi(y,y^{n},\psi)\geq\beta\}$ is closed, where $\beta\equiv\beta(\alpha)$ was defined in (8). Indeed, it is a super-level set of $\pi(\cdot,y^{n},\psi)$, and the latter is upper semicontinuous. This is because $\{y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n+1}_{-(n+1)},y_{n+1})\}$ is a closed set, so $\mathbbm{1}[\{y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n+1}_{-(n+1)},y_{n+1})\}]$ is upper semicontinuous, and the latter is preserved by finite averages. Thus, $\kappa(\alpha,y^{n},\psi)$ is a closed subset of the compact space $Y$, so it is itself compact.
Weak measurability of $\kappa$. Let $\Lambda\coloneqq[0,1]\setminus S_{n+1}$ and fix a non-empty open set $G\subseteq Y$. Because $Y$ is compact metric, it admits a countable dense set $Q\subseteq Y$; put $D_{G}\coloneqq Q\cap G$. Define
$f_{G}(\alpha,y^{n},\psi)\coloneqq\sup_{y\in D_{G}}\pi(y,y^{n},\psi),\qquad(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}.$
Step 1: each coordinate map is Borel. Fix $y\in D_{G}$. For $i\in\{1,\dots,n+1\}$, let $g_{i}(y^{n},\psi)\coloneqq\psi(y^{n+1}_{-i},y_{i})-\psi(y^{n},y),$ which is continuous in $(y^{n},\psi)$ because evaluation of $\psi$ in the sup-norm is continuous and coordinate deletion is continuous. Hence, the set $\{(y^{n},\psi):g_{i}\geq 0\}$ is closed, its indicator is Borel, and
$(y^{n},\psi)\mapsto\pi(y,y^{n},\psi)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}[(y^{n},\psi):g_{i}\geq 0]$
is a finite sum of Borel functions, so Borel.
Step 2: the countable supremum is Borel. Because $D_{G}$ is countable, $f_{G}$ is the pointwise supremum of a countable family of Borel functions; Therefore $f_{G}$ itself is Borel.
Step 3: translate the selector set. Since $\pi$ takes values in the finite set $S_{n+1}$, define $\beta(\alpha)\coloneqq\min\{s\in S_{n+1}:s>\alpha\},$ a piecewise-constant (hence Borel) map $\beta:\Lambda\to S_{n+1}$. Then
$\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\;\iff\;f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha).$
Consequently the selector set
$\Theta_{G}\coloneqq\bigl{\{}(\alpha,y^{n},\psi):f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha)\bigr{\}}$
is Borel in $\Lambda\times Y^{n}\times\mathscr{F}$. Thus
$(\kappa\cap G)^{-1}(\neq\emptyset)=\Theta_{G}\in\mathcal{B}\bigl{(}\Lambda\times Y^{n}\times\mathscr{F}\bigr{)},$
where $(\kappa\cap G)^{-1}(\neq\emptyset)\coloneqq\{(\alpha,y^{n},\psi):\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\}$, proving that $\kappa$ is weakly measurable.
Weak measurability of $\mathsf{CRED}$. Let $G\subseteq\Delta_{Y}$ be open. Because $\Delta_{Y}$ is metric and $\mathsf{CRED}$ has compact values, $\mathsf{CRED}(y^{n},\psi)\cap G\neq\emptyset$ if and only if there exists $P\in G$ satisfying all the constraints $P(A_{j})\leq\overline{\Pi}_{(y^{n},\psi)}(A_{j})$, where the $A_{j}$’s are the closed sets in $\Sigma_{Y}$. For each integer $m$, let $G_{m}=\{P\in G:d(P,G^{c})\geq 1/m\}$; $\{G_{m}\}$ is a countable basis for $G$. Hence
$\bigl{\{}(y^{n},\psi):\mathsf{CRED}(y^{n},\psi)\cap G\neq\emptyset\bigr{\}}\;=\;\bigcup_{m=1}^{\infty}\bigcup_{P\in\mathscr{D}_{m}}\bigl{\{}(y^{n},\psi):P(A_{j})\leq\overline{\Pi}_{(y^{n},\psi)}(A_{j}),\;\forall j\bigr{\}},$
where $\mathscr{D}_{m}$ is a countable dense subset of $G_{m}$ (using separability of $\Delta_{Y}$). Each inner set is Borel by Lemma 4; Countable unions preserve measurability, so the whole pre-image is in $\mathcal{B}\bigl{(}Y^{n}\!\times\!\mathscr{F}\bigr{)}$.
Weak measurability of $\mathsf{IHDR}$. For open $G\subseteq Y$, let =$\mathscr{B}=\{B_{k}\}_{k\in\mathbb{N}}$ be a countable closed basis of $Y$. Define $\phi_{k}(\alpha,\mathcal{M})=1-\alpha-\underline{P}_{\!\mathcal{M}}(B_{k}).$ Each $\phi_{k}$ is u.s.c. (by (iv), since $\mathscr{C}$ is endowed with the Vietoris topology and $\mathcal{M}$ is weak-compact), hence Borel on $[0,1]\times\mathscr{C}$. Because $\mathsf{IHDR}(\alpha,\mathcal{M})\cap G\neq\emptyset\iff\exists k:B_{k}\subseteq G,\;\phi_{k}(\alpha,\mathcal{M})\leq 0,$ the desired pre-image is a countable union of Borel sets, hence Borel.
We have now showed that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are uniformly compact-valued and weakly measurable, and thus morphisms of $\mathbf{WMeas}_{\text{uc}}$. Fix now any $\alpha\in[0,1]\setminus S_{n+1}$. It is easy to see that the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ remain uniformly compact-valued and weakly measurable. In addition, as in the Proof of Theorem 5, we have that $t(\mathsf{CRED})=s(\mathsf{IHDR}_{\alpha})=\mathscr{C}$.
Now, recall that the product of two Polish spaces is Polish. Since $Y$ is compact metric by (i’), it is also measurable Polish; $[0,1]\times\mathscr{C}$ is measurable Polish in the product topology by (iv); $Y^{n}\times\mathscr{F}$ is measurable Polish in the product topology by (i’) and (ii’); $[0,1]\times Y^{n}\times\mathscr{F}$ is measurable Polish in the product topology by (i’) and (ii’). Hence, these spaces are all objects of $\mathbf{WMeas}_{\text{uc}}$.
The same considerations that we put forth in Section 3 on the importance of Theorem 5, still hold here for Theorem 8.
@@
Theorems 5 and 8 tell us that the Full Conformal Prediction Diagram commutes when we focus either on the continuity or on the measurability aspects of the conformal prediction correspondence $\kappa$. This further testifies that, under our assumptions, uncertainty quantification is a built-in feature of Conformal Prediction methods.
The following remark shows that the conditions in Theorem 8 are truly minimal.
Remark 2 (Minimality of Hypotheses (i’), (ii’), (iii), (iv)).
  • Drop (ii’) (Polishness of $\mathscr{F}$).
[…]
    (b) Continuity. If continuity is also abandoned (e.g. $\psi(y^{1},y)=\mathbbm{1}[y\in\mathbb{Q}]$), $\kappa$’s graph is not Borel, so weak measurability fails.
  • Drop (iii) (no-tie rule). Pick $\alpha=k/(n+1)$, $k\in\{1,\ldots,n+1\}$. Then, $\kappa(\alpha,y^{n},\psi)$ is open but not closed, so individual values are not compact, and uniform compactness is lost.
  • Drop (iv) (weak/Vietoris topologies).
    (i) Put Total Variation on $\Delta_{Y}$. Evaluation is still 1-Lipschitz $\bigl{(}|P(A)-Q(A)|\leq d_{\text{TV}}(P,Q)\bigr{)}$, but $\Delta_{Y}$ is not compact in TV. Hence the images $\mathsf{CRED}(y^{n},\psi)$ need not be compact (take the sequence of Dirac masses $\delta_{1/k}$ inside $\mathsf{CRED}$); Uniformly compact-valuedness fails.
    (ii) Replace Vietoris with the Hausdorff metric on $\mathscr{C}$. Credal-set limits can break lower-probability continuity; $\mathsf{IHDR}$ fails to be weakly measurable, see Remark 1.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). Hence (i’), (ii’), (iii), (iv), together with consonance, are jointly minimal for the uniform compact-valuedness and weak measurability of $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$.

### v3

@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
Theorem 5 clarifies the nature of Conformal Prediction: Although often billed as an uncertainty quantification tool, CP furnishes a representation of predictive uncertainty, but not a quantification in the cardinal sense. The distinction is subtle yet essential: CP Regions depend only on the ranks of the non-conformity scores: For any strictly increasing transform $f$, $\kappa_{\alpha}(\psi)=\kappa_{\alpha}(f\circ\psi)$. Thus, while $\mathrm{diam}(\mathscr{R})$ is numerically well-defined once a metric on $Y$ is fixed, [^7] A similar argument holds if we consider the volume in place of the diameter of the CPR. there is no canonical cardinal scale of “amount of uncertainty” attached to $\mathscr{R}$. In particular,
$\mathrm{diam}(\mathscr{R}_{1})=3\mathrm{diam}(\mathscr{R}_{2})\not\Rightarrow\text{``$\mathscr{R}_{1}$ is three times more uncertain than $\mathscr{R}_{2}$''.}$
CP natively supports ordinal comparisons (larger/smaller regions), not cardinal arithmetic.
The commuting factorization $\kappa_{\alpha}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$, however, reveals a hidden layer. Indeed, computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ is a credal set (i.e. it “lives” in the space of probability measures), existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024); Chau et al., (2025)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
@@
    Dropping (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with its usual (non-compact) topology, take $n=1$, and define a continuous, permutation-invariant score that depends only on the first argument,
    $\psi(y^{1},y)\coloneqq g(y^{1}),\qquad g(x)=\begin{cases}1-e^{x},&x\leq 0,\\ x,&0<x<1,\\ 1,&x\geq 1.\end{cases}$
    Fix the closed set $A\coloneqq(-\infty,0]$ and write $x$ for $y^{1}$. For $n=1$ the conformal transducer satisfies $\pi(y,x)=\tfrac{1}{2}\bigl{(}1+\mathbbm{1}[g(y)\geq g(x)]\bigr{)}.$ Hence
    $\overline{\Pi}_{x}(A)=\sup_{y\leq 0}\pi(y,x)=\begin{cases}1,&g(x)<\sup_{y\leq 0}g(y)=1,\\ \tfrac{1}{2},&g(x)\geq 1.\end{cases}$
    Since $\sup_{y\leq 0}g(y)=1$ is not attained on $A$ while $g(1)=1$, we obtain $\overline{\Pi}_{x}(A)=\mathbbm{1}[x<1]+\tfrac{1}{2}\mathbbm{1}[x\geq 1],$ which has an upward limsup at $x=1$, $\limsup_{x\uparrow 1}\overline{\Pi}_{x}(A)=1>\tfrac{1}{2}=\overline{\Pi}_{1}(A).$ Thus the envelope is not upper semicontinuous at $x=1$. Consequently, $\mathsf{CRED}$ fails to be u.h.c. without compactness of $Y$.
@@
    Dropping (ii) (joint continuity of $\psi$). Let $Y=[0,1]$ (compact Hausdorff) and $n=1$. Define the permutation-invariant but discontinuous score
    $\psi(y^{1},y)\coloneqq g(y^{1}),\qquad g(x)=\mathbbm{1}[x\in\mathbb{Q}].$
    Fix $\alpha=\tfrac{3}{4}$ so that $\beta(\alpha)=1$, and take $y^{1}=x$. Then for $n=1$, $\pi(y,x)=\tfrac{1}{2}\bigl{(}1+\mathbbm{1}[g(y)\geq g(x)]\bigr{)}.$ If $x\in\mathbb{Q}$, $g(x)=1$ and
    $\kappa(\alpha,x,\psi)=\{y\in[0,1]:g(y)\geq 1\}=[0,1]\cap\mathbb{Q}.$
    If $x\notin\mathbb{Q}$, then $g(x)=0$ and $\pi(y,x)=1$ for all $y$, hence $\kappa(\alpha,x,\psi)=[0,1].$ Let $t\in(0,1)\setminus\mathbb{Q}$ be fixed and set $V\coloneqq[0,1]\setminus\{t\}$, which is open in the subspace topology and satisfies $\kappa(\alpha,x,\psi)\subset V$ when $x\in\mathbb{Q}$. For any neighborhood $U$ of a rational $x$, there exists $x^{\prime}\in U\setminus\mathbb{Q}$ with $\kappa(\alpha,x^{\prime},\psi)=[0,1]\not\subset V$. Thus $\kappa$ is not upper hemicontinuous at rational $x$. Joint continuity of $\psi$ is therefore necessary.
@@
    Dropping (iii) (no-tie rule). Let $\psi\in\mathscr{F}$ be the constant score $\psi(y^{n},y)\equiv 0$. Then for every $y,y^{n}$, all nonconformity scores are equal, hence $\pi(y,y^{n})=1$. Take $\alpha=1\in S_{n+1}$ (violating (iii)). Because $\kappa$ uses the strict inequality $\pi>\alpha$,
    $\kappa(1,y^{n},\psi)=\emptyset.$
    For any sequence $\alpha_{k}\uparrow 1$ with $\alpha_{k}<1$, we have $\kappa(\alpha_{k},y^{n},\psi)=Y$ for all $k$. Therefore $\kappa$ is not upper hemicontinuous at $\alpha=1$; The no-tie parameter restriction is essential.
@@
    Dropping (iv) (topology on $\mathscr{C}$). Endow $\Delta_{Y}$ with the weak topology, but equip $\mathscr{C}$ with the lower Vietoris (hit) topology, generated by the subbasic sets $\langle U\rangle_{-}\coloneqq\{\mathcal{M}\in\mathscr{C}:\mathcal{M}\cap U\neq\emptyset\}$, where $U\subset\Delta_{Y}$ is weak open. Let $Y=[0,1]$ (so (i) holds) and fix $c\in(0,1]$. Set $A\coloneqq\{0\}$, $P_{0}\coloneqq\delta_{0}$, $P_{c}\coloneqq\delta_{c}$, and define
    $\mathcal{M}_{k}\coloneqq\operatorname{co}(\{P_{0},P_{c}\})\in\mathscr{C}\quad(k\geq 1),\qquad\mathcal{M}_{\infty}\coloneqq\{P_{0}\}\in\mathscr{C},$
    where $\text{co}(\cdot)$ denotes the convex hull operator. Then, $\mathcal{M}_{k}\to\mathcal{M}_{\infty}$ in the lower Vietoris topology: For every weak open neighborhood $U$ of $P_{0}$ we have $P_{0}\in\mathcal{M}_{k}\cap U$ for all $k$. For any $E\in\Sigma_{Y}$,
    $\underline{P}_{\mathcal{M}_{k}}(E)=\inf_{P\in\mathcal{M}_{k}}P(E)=\min\{P_{0}(E),P_{c}(E)\},\qquad\underline{P}_{\mathcal{M}_{\infty}}(E)=P_{0}(E).$
    Fix any $\alpha^{\star}\in(0,1)\setminus S_{n+1}$, so $1-\alpha^{\star}\in(0,1)$. Then
    $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{\infty})=\bigcap\{E\in\Sigma_{Y}:P_{0}(E)\geq 1-\alpha^{\star}\}=\{0\},$
    whereas
    $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{k})=\bigcap\{E\in\Sigma_{Y}:\min(1_{\{0\}\subset E},1_{\{c\}\subset E})\geq 1-\alpha^{\star}\}=\{0,c\}\text{, }k\geq 1.$
    Let $V$ be any open neighborhood of $\{0\}$ with $c\notin V$ (e.g. $V=[0,\varepsilon)$ with $0<\varepsilon<c$). Then $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{\infty})\subset V$, but $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{k})=\{0,c\}\nsubseteq V$ for all $k$. Therefore $\mathsf{IHDR}_{\alpha^{\star}}$ is not upper hemicontinuous at $\mathcal{M}_{\infty}$ when $\mathscr{C}$ is endowed with the lower Vietoris topology. Hence the (full) Vietoris topology in (iv) is indispensable.
@@
An in-depth study of some of the properties of $\mathbf{UHCont}$ is carried out in Section 6. Now, instead, we briefly introduce structure $\mathbf{WMeas}_{\text{uc}}$, which will be explored in depth in Appendix A. There, we prove that it is indeed a category, and that the Full CP Diagram commutes in $\mathbf{WMeas}_{\text{uc}}$ as well, under essentially the same conditions of Theorem 5.
@@
  • Objects: Compact Polish spaces $(X,\Sigma_{X})$, with $\Sigma_{X}=\mathcal{B}(X)$, the Borel $\sigma$-algebra.
@@
  • Morphisms: Weakly measurable (see Definition 26 in Appendix C), uniformly compact-valued (see Definition 27 in Appendix C) correspondences $\Phi:X\rightrightarrows Y$.
@@
The Full Conformal Prediction Diagram commuting when we focus either on the continuity or on the measurability aspects of the conformal prediction correspondence $\kappa$ further testifies that, under our assumptions, uncertainty quantification is a built-in feature of Conformal Prediction methods.